80 likes | 181 Vues
This report reviews the Tevatron's beam physics progress from August 2004 to February 2005. During this period, the luminosity reached an average peak of 92.9, a 10% increase from the end of January. Challenges included vacuum valve issues and separator sparks, but notable improvements in beam lifetimes and reliability were achieved. Efforts to optimize tune and emittance led to successful operational diagnostics. Future goals focus on increasing proton density and enhancing luminosity stability while addressing ongoing system upgrades and personnel changes.
E N D
Tevatron progress: Aug’04 to Feb’05 • Luminosity: (average initial peak L) • Aug.04 (7 stores #3657-3678) L=84.4 • End of Jan (9 stores #3925-3953) L=92.9 (+10%) • -6% N_p, +6% N_pbar, (*eff ) reduced -10% • record L as of Aug.04 103e30 store #3657 record after FY04 shutdown 105.2e30 store #3952 • ~Same lifetimes: L 8.18.7 hrs, eff1617 hrsN_pbar 2522hrs, N_p ~90hrs • Int L is about same (15-18.5 then 16 pb-1/wk now) • FY’05 goal integral 470 pb-1 (now ~100 pb-1)
Shutdowns • FY’04 shutdown : 12 weeks, 28 jobs, including • SNEG pipes at D0 and CDF LB Q columns BH, McGee • 2 separators @D17+polarity switches RM et al • 412 dipoles reshimmed D.Harding • 7 um FWs at E11 N.Eddy et al • Kaiser coil and roll measmnts JA, J.Volk • B0 and D0 HLS systems installed J.Volk et al • 4 houses warmed up MechSupport • Short 3 days shutdown Feb’05: 15 jobs including • F17 kicker magnet replaced Bruce • Crystal collimator fixed Dean • Vacuum leak A49 fixed Scott
Highlights on Operations/Reliability: • Beam startup took 24 shifts ( ~ as planned) • A bit slowed by troubles with vacuum valves Dean! • Reached pre-shutdown L~60e30 after 9 stores • Suffered from separator sparks for ~12 days in Dec’04 • 3 stores lost , 2 damaged • lowered voltages by 10% on Dec 18, 2004 one spark since then • Mixed source shots regular very long stores (record 45.3hrs in #3920) • Stores lost due to QFA4, cryo(1), controls(1), RF trips (5 trips, 1 lost), longitudinal damper off (1) • D0 and (less) CDF halo rates are down, LOSTP spiky after 20hrs, no motion at 12 Hz • Superb lifetimes at 150, ramp losses low • Losses in squeeze 1-2% p/pbars, dropped Q’ 208, still… • Scallops – were regular, disappeared after dQ -0.002 change • No transv-dampers at injection either higher Q’ or Octupoles • CDF and D0 IP positions and angles corrected (upto 3mm) –with correctors or by alignment of B0 LB quadrupoles • Lifetime at 150 is great, losses on ramp smaller • P-lifetime at LB wandering 60hrs160hrs60hrs…
Tevatron Beam Physics Progress: • Octupoles: O1 O2 commisioned for differential chromaticity control, OZF OZD for stability at 150 chromaticit reduced as well as losses (esp pbar) (Petr, Yuri and Jerry) • Longitudinal blowup is multibunch mode 1 (Tan) • Lattice re-measured after shutdown at 150 & LB = OK (Valery) • WP scan at EoS better proton tunes (ZXL, VS, JPC) • Pbar bunch-by-bunch orbits at 150 as modeled (Sasha Valishev) • Tev B-Beam Phenomena paper in progress (many) • Wake field effects seen in head-tail motion (Vahid) • Beam-beam modeling Q’=0 and 23 RFC idea (Velery, Sasha Valishev) • Theory of b-b resonance diffusion with noise (Yuri) • P-loss rate analysis at HEP and 150 ~ Na/emma^2at (VS) • IBS emittance growth study, 2100 hrs lifetime (Vladimir, Alvin) • Analysis of losses of pbars from RR and AA (Aimin, Tanaji) • MP9 studies of separator sparking 10 times /10kV (Prokofiev)
Diagnostics: • Abort gap monitor operational (Randy, Eugene) • New BPMs at A3 and B3, see pbars, TBT (SW, JS, BW, et al), finish in April’05 • Noisy MUX removed better 1.7GHz spectra (Andreas, Ralph) • 7 um FWs @E11 acceptable for CDF&D0 (Nathan Eddy) • SBD reports dP/P and long-Emittance (Bob Flora) • First results from long phase detector (Aisha, BW, JPC) • BBQ shows 60 Hz lines here and there (Tan) • HLS systems at B0 ~works (Jim Volk) • 90% of p-SLite found, nice modeling, progress toward more believable data (Randy) • Q’ from 21MHz Schottky under test (TSen, Paul) • IPM electronics to be ready at the end of March (AJ et al)
Misc: • MTF measurement mini-review : 0.7 unit b1 drift, broken anchors not a problem (Pierre B. et al) • BBCompensation review –OK’d FY’05 plan (VS, ZXL, et al) • CDF/D0 lumi ratio still unresolved, D0 admitted errors (Vaia et al) • BTeV cancelled • 6 invited speakers for PAC’05 from Tevatron • Personnel: Out Aimin Xiao AID Claudio Rivetta Vinay Boocha Sasha Valishev Kip Bishofberger (Mar) • In: Jean-Paul Carneiro (Jan) Elmie Peoples (Jan) Seva Kamerdzhiev (~end of March) Eliana Wendt (~mid March) Sasha Valishev (~April) Alex Koschik (offer sent) • Part-timers MM Run II Upgr, TSLARP (Tan+Andreas active) • Letter of recognition to Mike McGee
Next Steps • Increase Np 230250280 (same Emm) 10…20% in L Tev coord • Lifetime improvements at HEP 6-12% in L • Octupoles@HEP to reduce chroma’s Petr/Yuri • Stabilize pbar tunes in stores Ron • Explore new proton WPs 7/12 < Q < 3/5 Jerry et al • Reduce * 35 28 cm 8-12% in L Sasha/Val • Reduce p-loss at 150 helix (octupoles, tunes) 2-4% in L coord • Fly only E11 Wires at 150 to save 1 pi 1-2% in L Nathan Other items: • Finish BPM installation and Commissioning (BPM project team) • Pbar SL mirror move in/out (Randy, Stephen, Todd) • Commission tune-tracker (CYTan) • Commission head-tail monitor and orbit stabilization (Vahid Ranjbar) • Test magnets, assemble and test TEL-2 (BBC team) • Install IPMs and OTRs (Andreas, Vic et al) • Minor SBD improvements, and FWs OAC with AT dP/P (Flora et al) • Commission Long Phase Monitor (Aisha, Jean-Paul) • Replace RF cavity water-heaters (JohnReid) • Make B0 & D0 HLS systems operational (Jim Volk) • First tests of Crystal Collimator (Dean)
Next Three Months: Expectations • Over the next 3 months, I expect: • stable operation at 90-100 e30, combined shots • D0 luminosity monitor fixed (+6% in AVG lumi) • less study time as reliability improves • possible improvements • 10-20% increase in proton intensity with ~same emittance • 5-10% more pbars from RR smaller Emm, 120mA stash • 5-10% lumi-lifetime improvements WPs, octupoles, etc • * 35 28 cm studies to start and may be even cashed in • As the result, I expect peak CDF luminosity to be around 120-130e30 by May 1, 2004